# Constrained Decoding for Secure Code Generation

### v1

@@ Abstract.
This paper introduces a new benchmark, CodeGuard+, along with two new metrics, secure-pass@$k$ and secure@$k_{\text{pass}}$, to measure Code LLMs’ ability to generate both secure and correct code. Using our new evaluation methods, we show that the state-of-the-art defense technique, prefix tuning, may not be as strong as previously believed, since it generates secure code but sacrifices functional correctness. We also demonstrate that different decoding methods significantly affect the security of Code LLMs.
@@
Furthermore, we explore a new defense direction: constrained decoding for secure code generation. We propose new constrained decoding techniques to generate code that satisfies security and correctness constraints simultaneously. Our results reveal that constrained decoding is more effective than prefix tuning to improve the security of Code LLMs, without requiring a specialized training dataset. Moreover, constrained decoding can be used together with prefix tuning to further improve the security of Code LLMs.
Large Language Models; Code Generation; Code LLM; Secure Code Generation; AI Safety
1. Introduction
@@
Code Large Language Models (Code LLMs) such as GitHub Copilot (GitHub, 2021) and Amazon CodeWhisperer (Amazon, 2023) have been used by millions of developers (Tiernan Ray, ZDNet, 2023). Research studies have shown that Code LLMs can significantly boost the productivity of developers (Eirini Kalliamvakou, GitHub Blog, 2022; Maxim Tabachnyk and Stoyan Nikolov, Google Research, 2022). However, Code LLMs are not secure: they may recommend code that contains security vulnerabilities. In particular, Pearce et al. have shown that 40% of programs generated by GitHub Copilot are vulnerable (Pearce et al., 2022). As developers increasingly rely on Code LLMs in their daily tasks, it is critical to ensure that LLM-generated code is secure.
@@
Prior works (Pearce et al., 2022; He and Vechev, 2023; Hajipour, Hossein and Hassler, Keno and Holz, Thorsten and Schönherr, Lea and Fritz, Mario, 2023; Wu et al., 2023) that automatically evaluate the security of code generated by LLMs focus on only security, while ignoring correctness. Correctness is an important criterion for developers to accept code suggested by LLMs. Thus, if a model generates secure but incorrect code, it is not meaningful for a developer. We argue that the previous evaluation method gives us a false sense of security when we compare different models; and this could overestimate the ability of defense techniques to generate secure code. As a result, this hinders the progress of the research community to build more secure Code LLMs.
@@
In this paper, we propose a new benchmark CodeGuard+ to evaluate the security of Code LLMs, and we study a new defense direction of using constrained decoding to enhance the security of Code LLMs. To propose new evaluation methods for Code LLMs, we face the following challenges. First, there is a disconnection between security evaluation and correctness evaluation. Existing benchmarks including HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2024), and MBPP (Austin et al., 2021) can evaluate correctness of Code LLMs, but they are not relevant to triggering security vulnerabilities such as command injection. On the other hand, security-relevant prompt datasets (Pearce et al., 2022; Siddiq and Santos, 2022) do not come with any test suite to evaluate correctness. To this end, we propose a new benchmark CodeGuard+. We modify the original security prompts from a widely used security prompt dataset (Pearce et al., 2022) to be suitable for tests, and we develop test cases to check correctness of code completions given these prompts.
@@
The second challenge is, the prior metric that evaluates security of Code LLMs overlooks functional correctness, which is not practical since developers prefer to accept correct code suggested by LLMs. Previous works calculate the security rate as the percentage of secure programs within unique generated programs that can be parsed and compiled (He and Vechev, 2023; Pearce et al., 2022). This does not measure correctness and forgives generated code that is functionally wrong. This is disconnected from the standard pass@$k$ metric (Chen et al., 2021) widely used in the literature for comparing performance of Code LLMs, which defines the expected likelihood of generating any correct code output within $k$ code outputs. Thus, we propose new evaluation metrics including secure-pass@$k$ and secure@$k_{\text{pass}}$. When $k=1$, the intuition is, secure-pass@$1$ measures the expected likelihood of generating both secure and semantically correct code given a single generation; and secure@$1_{\text{pass}}$ measures the likelihood of any generated correct code being secure.
@@
Furthermore, we study a new defense direction of constrained decoding for secure code generation. In actuality, a pre-trained Code LLM does not give us a mapping from an input to an output, but instead, it models the conditional probability distribution of outputs given a prompt. To generate a concrete output from a Code LLM, a decoding procedure is used to search over the output space using conditional probability. Prior works in this space consider the decoding procedure as a black-box function. In this paper, we open up the black box and demonstrate new opportunities to improve the security of Code LLMs. We formulate a new constrained decoding problem to generate secure and correct code. This problem is given a set of constraints to enforce correctness and security for the generated program. Then, given a prompt and a pre-trained Code LLM, the constrained decoding task needs to generate code that satisfy all the specified constraints.
We specify correctness constraints and security constraints for code generated by prompts in our benchmark CodeGuard+. We formulate correctness constraints based on our understanding of the prompts and the anticipated behaviors of the resulting code. To specify security constraints, we use knowledge about common secure coding practices, semantics of the prompts, and the corresponding vulnerability type (CWE) that might be triggered by the prompt. Here are some examples of common secure coding practices. To avoid out-of-bound write, we need the generated code to do the array index bound check for the allocated buffer. To process untrusted user input, the generated code should perform input validation. Even though writing specifications is a manual process, having security domain knowledge from an undergraduate-level security class is sufficient to specify constraints.
Next, we propose two techniques to enforce our constraints, in two kinds of decoding methods, respectively: autoregressive and non-autoregressive decoding. Autoregressive decoding generates output tokens one at a time, in a left-to-right manner. We find that sampling-based methods work better than deterministic methods to generate secure code if we do autoregressive decoding. At every step of decoding, a deterministic method always has one best output, which has a high risk of eventually leading to vulnerable code. Whereas, a sampling-based method has more opportunities for exploration. Therefore, we propose a Constrained Beam Sampling technique to enforce our constraints while avoiding the pitfalls of being stuck in vulnerable code solutions during the generation.
@@
Figure 1. We compare CodeGen + Prefix-tuning model, trained by the state-of-the-art defense (He and Vechev, 2023), against the baseline CodeGen model. Our new metric secure-pass@$1$ is more realistic than SVEN Security Rate used in (He and Vechev, 2023), since we evaluate both security and correctness of generated code, while SVEN Security Rate does not evaluate correctness. SVEN Security Rate severely overestimates how secure a model really is. Under Beam Sampling, secure-pass@$1$ of CodeGen + Prefix-tuning is no longer better than CodeGen.
@@
We propose a second constrained decoding technique by adapting a gradient-based non-autoregressive decoding method, MuCoLa (Kumar et al., 2022). Non-autoregressive decoding generates all tokens in the output altogether, instead of one token at a time. These methods are gradient-based. They start by initializing all the tokens in the output sequence, and then iteratively update the tokens using gradients of some function, e.g., language model loss function. In the non-autoregressive generation paradigm, MuCoLa is a state-of-the-art technique for constrained text generation. It formulates decoding as sampling from an energy-based model using Langevin Dynamics. To adapt MuCoLa for secure code generation, we define our own energy function that is more suitable to enforce our constraints.
@@
Using our benchmark CodeGuard+ and new evaluation metrics, we thoroughly evaluate a baseline model against a secure model trained with the state-of-the-art prefix tuning defense. In particular, we run Nucleus Sampling and Beam Sampling over two models, CodeGen-2.7B as the baseline, and CodeGen-2.7B trained using the prefix tuning method SVEN (He and Vechev, 2023). Figure 1 highlights some results. Using Nucleus Sampling, CodeGen + Prefix-tuning model has an 84.49% SVEN security rate, 23.84% higher than the baseline. However, since SVEN security rate does not measure correctness, this severely overestimate how secure CodeGen + Prefix-tuning model really is. When we use our new metric to measure both security and correctness of generated code, CodeGen + Prefix-tuning model has only 47.48% secure-pass@$1$, almost half of the original security rate, and only 4.31% better than secure-pass@$1$ of the baseline CodeGen. We observe that prefix tuning sacrifices functional correctness to generate secure code, as CodeGen + Prefix-tuning model has only 55.6% pass@$1$. When using Beam Sampling, the performance of both models becomes better than using Nucleus Sampling. However, in this case, prefix tuning no longer has any advantage over the baseline model when evaluated using secure-pass@$1$. Our results indicate that the state-of-the-art defense may not be as strong as previously believed.
@@
Last but not least, we evaluate our new constrained decoding schemes over CodeGen-2.7B, CodeGen-2.7B with prefix tuning, and StarCoder2-3B. Figure 2 highlights some key results. Overall, constrained decoding performs better than unconstrained decoding. In particular, constrained decoding over CodeGen (76% secure-pass@$1$) works better than prefix tuning with unconstrained decoding (62.19% secure-pass@$1$). The advantage of decoding is, it does not require specialized training datasets as needed by prefix tuning (He and Vechev, 2023) and instruction tuning (He et al., 2024). We also show that constrained decoding can be used together with prefix tuning to further improve the security of Code LLMs. Figure 2 shows that constrained decoding further improves the performance of CodeGen + Prefix-tuning model from 62.19% to 82.17% secure-pass@$1$.
Figure 2. Our constrained decoding technique can improve secure-pass@$1$ for both the baseline CodeGen model and the defended CodeGen + Prefix-tuning model (He and Vechev, 2023), compared to unconstrained decoding. Constrained decoding over CodeGen has 13.81% higher secure-pass@$1$ than CodeGen + Prefix-tuning with unconstrained decoding. Constrained decoding can be used with the prefix tuning defense together.
@@
Our CodeGuard+ and source code are available at https://github.com/Dynamite321/CodeGuardPlus. Our contributions are summarized as follows:
@@
  • We study a new defense direction of using constrained decoding to generate secure code. We formulate the problem, propose correctness and security constraints, and we propose two constrained decoding techniques.
@@
  • We evaluate our constrained decoding techniques over CodeGen and StarCoder2. We show that constrained decoding can significantly improve the security of Code LLMs. Constrained decoding can be used together with prefix tuning defense to further boost the performance.
2. Background and Related Work
@@
Large tech companies have developed closed-source Code LLMs such as GitHub Copilot (GitHub, 2021), Amazon CodeWhisperer (Amazon, 2023), Google’s PaLM (Chowdhery et al., 2023), and those with paid API services from OpenAI and Anthropic. On the other hand, several communities have released open-source Code LLMs. To rank the quality of Code LLMs, it is standard to use the pass@$k$ metric (Chen et al., 2021) over benchmark datasets such as HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2024) and MBPP (Austin et al., 2021). The pass@$k$ metric represents the likelihood of any one out of k generations passing the unit tests when evaluated over a dataset. In our work, we experiment with open-source Code LLMs as they have demonstrated competitive performance as closed-source ones. Specifically, we experiment with CodeGen (Nijkamp et al., 2022) and StarCoder2 (Lozhkov et al., 2024). CodeGen (Nijkamp et al., 2022) is pre-trained using next-token prediction language modeling, and three open-source code datasets supporting six programming languages. StarCoder2 (Lozhkov et al., 2024) is pre-trained using the fill-in-the-middle task, over both code and text from GitHub, and other natural language datasets.
@@
Since Code LLMs are trained with source code written by developers, they have learned vulnerable code patterns from humans. Pearce et al. (Pearce et al., 2022) show that 40% of programs generated by GitHub Copilot are vulnerable. Similar results are supported by another study (Fu et al., 2024). Researchers have used different prompting techniques for Code LLMs to generate vulnerable source code. For example, zero-shot prompting (Khoury et al., 2023; Tihanyi et al., 2023), few-shot prompting (Hajipour, Hossein and Hassler, Keno and Holz, Thorsten and Schönherr, Lea and Fritz, Mario, 2023), prompt tuning using natural language (Wu et al., 2023), mining prompts from StackOverflow (Hamer et al., 2024), and using developer-written code preceding vulnerable code (Bhatt et al., 2023). Elgedawy et al. (Elgedawy et al., 2024) wrote 9 new tasks to prompt ChatGPT, BARD, and Gemini to generate code, used ground rules to check the functional correctness of outputs, and manually checked the security of the outputs. Previously, there was no automated evaluation to check both correctness and security.
User studies have shown that developers who have access to AI coding assistants backed by Code LLMs do not write more insecure code if they write in low-level C language (Sandoval et al., 2023); but they write significantly less secure code if they write in Python or JavaScript, to do encryption/decryption, sign messages, or process untrusted input from users (Perry et al., 2023).
@@
Recently, researchers have used prompt engineering (Homoliak et al., 2024), prefix tuning (He et al., 2024), instruction tuning (He et al., 2024), and vulnerability repair (Pearce et al., 2023) to help Code LLMs generate secure code. Notably, prefix tuning (He and Vechev, 2023) has achieved promising results. Prefix is a sequence of continuous vectors, prepended to the input (Li and Liang, 2021). The trainable parameters in the prefix should capture task-specific information, i.e., the task to generate secure code or vulnerable code. Prefix tuning only needs to train 0.1% of parameters in a model, which is more lightweight than instruction tuning that trains all model parameters. Using prefix tuning, He and Vechev (He and Vechev, 2023) can increase the ratio of secure code in programs generated by CodeGen-2.7B from 59% to 92%. Given vulnerable code, researchers have explored vulnerability repair using reinforcement learning with LLMs (Islam and Najafirad, 2024) and zero-shot repair using LLMs (Pearce et al., 2023). Pearce et al. (Pearce et al., 2023) suggest that it is challenging to maintain correctness in repaired code.
@@
Constrained decoding methods have been proposed for text generation, such as generalizing image captioning to out-of-distribution scenes (Anderson et al., 2017), enforcing lexical constraints in neural machine translation (Post and Vilar, 2018), and including common sense in outputs (Lu et al., 2021; Kumar et al., 2022, 2021). The benefit of constrained decoding is that we do not need any training or fine-tuning over a pre-trained model. Decoding methods for code generation have not received much attention. Storhaug et al. (Storhaug et al., 2023) have experimented with blocking vulnerability-related keywords when generating smart-contract code, but they have not experimented with positive constraints. To the best of our knowledge, we are the first to study the performance of different decoding algorithms for secure code generation.
3. New Evaluation Guidelines

### v2

@@ Abstract
This paper introduces a new benchmark, CodeGuard+, along with two new metrics, to measure Code LLMs’ ability to generate both secure and correct code. Using our new evaluation methods, we show that the state-of-the-art defense technique, prefix tuning, may not be as strong as previously believed, since it generates secure code but sacrifices functional correctness. We also demonstrate that different decoding methods significantly affect the security of Code LLMs.
@@
Furthermore, we explore a new defense direction: constrained decoding for secure code generation. We propose new constrained decoding techniques to generate secure code. Our results reveal that constrained decoding is more effective than prefix tuning to improve the security of Code LLMs, without requiring a specialized training dataset. Moreover, our evaluations over eight state-of-the-art Code LLMs show that constrained decoding has strong performance to improve the security of Code LLMs, and our technique outperforms GPT-4.
Keywords:
Large Language Models, Code Generation, Code LLM, Secure Code Generation, AI Safety.
1 Introduction
@@
Code Large Language Models (Code LLMs) such as GitHub Copilot [1] and Amazon CodeWhisperer [2] have been used by millions of developers [3]. Research studies have shown that Code LLMs can significantly boost the productivity of developers [4, 5]. However, Code LLMs are not secure: they may recommend code that contains security vulnerabilities. In particular, Pearce et al. have shown that 40% of programs generated by GitHub Copilot are vulnerable [6]. As developers increasingly rely on Code LLMs in their daily tasks, it is critical to ensure that LLM-generated code is secure.
@@
Prior works [6, 7, 8, 9] that automatically evaluate the security of code generated by LLMs focus on only security, while ignoring correctness. Correctness is an important criterion for developers to accept code suggested by LLMs. Thus, if a model generates secure but incorrect code, it is not meaningful for a developer. We argue that the previous evaluation method gives us a false sense of security when we compare different models. This could overestimate the ability of defense techniques to generate secure code. As a result, this hinders the progress of the research community to build more secure Code LLMs.
@@
In this paper, we propose a new benchmark CodeGuard+ to evaluate the security of Code LLMs, and we study a new defense direction of using constrained decoding to enhance the security of Code LLMs. To propose new evaluation methods for Code LLMs, we face the following challenges. First, there is a disconnection between benchmarks for security evaluation and correctness evaluation. Existing benchmarks including HumanEval [10], HumanEval+ [11], and MBPP [12] can evaluate correctness of Code LLMs, but they are not relevant to triggering security vulnerabilities such as command injection. On the other hand, security prompt datasets [6, 13] do not come with any test suite to evaluate correctness. To this end, we propose a new benchmark CodeGuard+. We modify the original prompts from previous security prompt datasets [6, 13] to be suitable for tests, and we develop test cases to check correctness of code completions given these prompts. Our benchmark has 91 prompts across 34 CWEs, larger than the state-of-the-art security prompt dataset that is widely used [6].
@@
The second challenge is that the prior metric that evaluates the security of Code LLMs overlooks functional correctness, which is not practical since developers prefer to accept correct code suggested by LLMs. Previous works calculate the security rate as the percentage of secure programs within unique generated programs that can be parsed and compiled [7, 6]. This does not measure correctness and forgives generated code that is functionally wrong. This is disconnected from the standard pass@$k$ metric [10] widely used in the literature for comparing performance of Code LLMs, which defines the expected likelihood of generating any correct code output within $k$ code outputs. Thus, we propose new evaluation metrics including secure-pass@$k$ and secure@$k_{\text{pass}}$. When $k=1$, the intuition is that secure-pass@$1$ measures the expected likelihood of generating both secure and semantically correct code given a single generation; secure@$1_{\text{pass}}$ measures the likelihood of any generated correct code being secure.
@@
Furthermore, we study a new defense direction of constrained decoding for secure code generation. In actuality, a pre-trained Code LLM does not give us a mapping from an input to an output, but instead, it models the conditional probability distribution of outputs given a prompt. To generate a concrete output from a Code LLM, a decoding procedure is used to search over the output space using the conditional probability distribution. Prior works in this space consider the decoding procedure as a black-box function. In this paper, we open up the black box and demonstrate new opportunities to improve the security of Code LLMs. We formulate a new constrained decoding problem to generate secure and correct code. This problem is given a set of constraints to enforce in the generated program. Then, given a prompt and a pre-trained Code LLM, the constrained decoding task needs to generate code that satisfies all the specified constraints.
@@
Figure 1: We compare CodeGen + Prefix-tuning model, trained by the state-of-the-art defense [7], against the baseline CodeGen model. Our metric secure-pass@$1$ is more realistic than SVEN Security Rate used in [7], since we evaluate both security and correctness of generated code, while SVEN Security Rate does not evaluate correctness. SVEN Security Rate severely overestimates how secure a model really is. The secure-pass@$1$ of CodeGen + Prefix-tuning is only 2.53% better than CodeGen with Beam Sampling.
We specify security constraints for code generated by prompts in our benchmark CodeGuard+. To specify the constraints, we use knowledge about common secure coding practices and the corresponding vulnerability type (CWE) that might be triggered by the prompt. For example, to avoid out-of-bound write, we need the generated code to do the array index bound check; to process untrusted user input, the generated code should perform input validation. Although writing specifications is a manual process, having security domain knowledge from an undergraduate-level security class is sufficient to specify constraints. All our constraints can be expressed as either a keyword or a template string, e.g., writing a function name, or filling out the variable name in the template for index bound check. Therefore, it is easy for developers to write constraints.
Next, we propose two techniques to enforce our constraints, in two kinds of decoding methods, respectively: autoregressive decoding and non-autoregressive decoding. Autoregressive decoding generates output tokens one at a time, in a left-to-right manner. We find that sampling-based methods work better than deterministic methods to generate secure code if we do autoregressive decoding. At every step of decoding, a deterministic method always has one output, which has a high risk of eventually leading to vulnerable code. Whereas, a sampling-based method has more opportunities for exploration. Therefore, we propose a Constrained Beam Sampling technique to enforce our constraints while avoiding the pitfalls of being stuck in vulnerable code solutions during the generation.
Figure 2: Our constrained decoding technique can improve secure-pass@$1$ of all six open-source Code LLMs of sizes ranging from 2.7B to 34B. Every model with constrained decoding shows better secure-pass@$1$ than GPT-4 with Nucleus Sampling.
@@
We propose a second constrained decoding technique by adapting a gradient-based non-autoregressive decoding method, MuCoLa [14]. Non-autoregressive decoding generates all tokens in the output altogether, instead of one token at a time. These methods are gradient-based. They start by initializing all the tokens in the output sequence, and then iteratively update the tokens using gradients of some function, e.g., language model loss function. In the non-autoregressive generation paradigm, MuCoLa is a state-of-the-art technique for constrained text generation. It formulates decoding as sampling from an energy-based model using Langevin Dynamics. To adapt MuCoLa for secure code generation, we define our own energy function that is more suitable to enforce our constraints.
Using our benchmark CodeGuard+ and new metrics, we thoroughly evaluate different decoding schemes over eight state-of-the-art Code LLMs with varied model sizes, including seven open-source models and one proprietary model GPT-4. The open-source models are: CodeGen-2.7B, SVEN (CodeGen-2.7B with prefix tuning), StarCoder2-3B, CodeGemma-7B, Llama3-8B, DeepseekCoder-33B, and CodeLlama-34B. Our results reveal that decoding methods make a big difference in generating secure and correct code, even without constraints. For six open-source models, Beam Sampling has higher secure-pass@$1$ than Nucleus Sampling, while the two methods have similar performance in only one open-source model.
@@
Our new metrics reveal a more realistic performance of the state-of-the-art prefix tuning defense. Figure 1 highlights some results. We run Nucleus Sampling and Beam Sampling over two models, CodeGen-2.7B as the baseline, and CodeGen-2.7B trained using the prefix tuning method SVEN [7]. Using Nucleus Sampling, CodeGen + Prefix-tuning has a 71.91% SVEN security rate, 18.26% higher than the baseline. However, since SVEN security rate does not measure correctness, this severely overestimates how secure CodeGen + Prefix-tuning really is. When we use our new metric for evaluation, CodeGen + Prefix-tuning has only 29.14% secure-pass@$1$, less than half of the original security rate, and only 3.07% better than secure-pass@$1$ of the baseline. We observe that prefix tuning sacrifices functional correctness to generate secure code, which decreases pass@$1$ by 6.94%. Our results indicate that the state-of-the-art defense may not be as strong as previously believed.
@@
Last but not least, we evaluate our new constrained decoding schemes over open-source Code LLMs. Our results show that constrained decoding over CodeGen (51.25% secure-pass@$1$) works better than prefix tuning with unconstrained decoding (36.3% secure-pass@$1$ for SVEN). The advantage of decoding is that it does not require specialized training datasets as needed by prefix tuning [7] and instruction tuning [15]. Figure 2 highlights that our Constrained Beam Sampling technique improves secure-pass@$1$ for all six open-source models of sizes ranging from 2.7B to 34B. Every model with constrained decoding outperforms GPT-4 with unconstrained decoding (Nucleus Sampling).
@@
Our CodeGuard+ benchmark is available at https://github.com/Dynamite321/CodeGuardPlus. Our contributions are summarized as follows:
@@
  • We study a new defense direction of using constrained decoding to generate secure code. We formulate the problem, propose security constraints, and we propose two constrained decoding techniques.
@@
  • We evaluate our constrained decoding techniques over eight state-of-the-art Code LLMs. We show that constrained decoding can significantly improve the security of Code LLMs. Our technique outperforms GPT-4.
2 Background and Related Work
@@
Large tech companies have developed closed-source Code LLMs such as GitHub Copilot [1], Amazon CodeWhisperer [2], Google’s PaLM [16], and those with paid API services from OpenAI and Anthropic. On the other hand, several communities have released open-source Code LLMs. To rank the quality of Code LLMs, it is standard to use the pass@$k$ metric [10] over benchmark datasets such as HumanEval [10], HumanEval+ [11] and MBPP [12]. The pass@$k$ metric represents the likelihood of any one out of k generations passing the unit tests when evaluated over a dataset. In our work, we experiment with state-of-the-art open-source Code LLMs as well as the proprietary GPT-4 [17]. The state-of-the-art open-source Code LLMs are typically pre-trained using a mix of text and source code datasets supporting multiple programming languages. For open-source Code LLMs, we experiment with CodeGen-2.7B [18], SVEN [7], StarCoder2-3B [19], CodeGemma-7B [20], Llama3-8B [21], DeepseekCoder-33B [22], and CodeLlama-34B [23].
@@
Since Code LLMs are trained with source code written by developers, they have learned vulnerable code patterns from humans. Pearce et al. [6] show that 40% of programs generated by GitHub Copilot are vulnerable. Similar results are supported by another study [24]. Researchers have used different prompting techniques for Code LLMs to generate vulnerable source code. For example, zero-shot prompting [25, 26], few-shot prompting [8], prompt tuning using natural language [9], mining prompts from StackOverflow [27], and using developer-written code preceding vulnerable code [28]. Elgedawy et al. [29] wrote 9 new tasks to prompt ChatGPT, BARD, and Gemini to generate code, used ground rules to check the functional correctness of outputs, and manually checked the security of the outputs. Previously, there was no automated evaluation to check both correctness and security.
User studies have shown that developers who have access to AI coding assistants backed by Code LLMs do not write more insecure code if they write in low-level C language [30]. However, they write significantly less secure code if they write in Python or JavaScript, to do encryption/decryption, sign messages, or process untrusted input from users [31].
@@
Recently, researchers have used prompt engineering [32], prefix tuning [7], instruction tuning [15], and vulnerability repair [33] to help Code LLMs generate secure code. Notably, prefix tuning [7] has achieved promising results. Prefix is a sequence of continuous vectors, prepended to the input [34]. The trainable parameters in the prefix should capture task-specific information, i.e., the task to generate secure code or vulnerable code. Prefix tuning only needs to train 0.1% of parameters in a model, which is more lightweight than instruction tuning that trains all model parameters. Using prefix tuning, He and Vechev [7] can increase the ratio of secure code in programs generated by CodeGen-2.7B from 59% to 92%. Given vulnerable code, researchers have explored vulnerability repair using reinforcement learning with LLMs [35] and zero-shot repair using LLMs [33]. Pearce et al. [33] suggest that it is challenging to maintain correctness in repaired code.
@@
Constrained decoding methods have been proposed for text generation, such as generalizing image captioning to out-of-distribution scenes [36], enforcing lexical constraints in neural machine translation [37], and including common sense in outputs [38, 14, 39]. The benefit of constrained decoding is that we do not need any training or fine-tuning over a pre-trained model. Decoding methods for code generation have not received much attention. Storhaug et al. [40] have experimented with blocking vulnerability-related keywords when generating smart-contract code, but they have not experimented with positive constraints. To the best of our knowledge, we are the first to study the performance of different decoding algorithms for secure code generation.
3 New Evaluation Guidelines